OpenStudy (anonymous):

Summation notation help? One problem. Will give medal! It says for each sum I need to find the number of terms, the first term, the last term, and to evaluate the sum. On top of the sigma is 8, under it is n=1, and to the right is 2n/3. I know the number of terms is 8. How do I find the rest.

OpenStudy (anonymous):

\[ \sum_{n=1}^{8}\frac{2n}3 = \frac{2(1)}{3} + \frac{2(2)}{3} + \frac{2(3)}{3} +\ldots + \frac{2(7)}{3} + \frac{2(8)}{3} \]

OpenStudy (anonymous):

\[ =(1+2+3+4+5+6+7+8)\times \frac{2}{3} \]

OpenStudy (anonymous):

36*(2/3)=24. That's the sum evaluated?

OpenStudy (anonymous):

@wio

OpenStudy (whpalmer4):

Yes, that is correct.

OpenStudy (whpalmer4):

\[\sum _{n=1}^8 k n=k \sum _{n=1}^8 n\]is the property of sums that @wio used

OpenStudy (whpalmer4):

and \[\sum _{n=1}^8 n=\frac{n(n+1)}{2} = \frac{8*9}{2} = 36\]

OpenStudy (anonymous):

Thanks so much @whpalmer4